### v1

# Jones’ conjecture for Halin graphs and a bit more

Abstract
We prove Jones’ famous conjecture for Halin graphs and a somewhat more general class of graphs too. A based planar graph is a planar one that has a face adjacent to every other face. We confirm Jones’ Conjecture for based planar graphs. Namely, if a based planar graph does not contain $k+1$ vertex-disjoint cycles, then it suffices to delete $2k$ vertices to make it acyclic.
Keywords: Jones’ conjecture, planar graph, cycle packing, feedback vertex set
1 Introduction
Jones’ conjecture [1] states that if a planar graph contains no more than $k$ pairwise disjoint cycles then there are $2k$ vertices such that removing them from the graph we get a forest. Jones’ conjecture was proved for outerplanar graphs [1] and subcubic planar graphs [2]. Currently, the best known general result is that if the maximum number of disjoint cycles in an arbitrary planar graph is $k$, then there are $3k$ vertices such that removing them from the graph we get a forest as proved independently by Chen et al. [3], Ma et al. [4] and Chappell et al. [5].
The main result of this paper is to prove the Jones’ conjecture for planar graphs that have a face that is adjacent to every other faces. Let’s call these graphs based planar graphs. (As a generalization of the based polyhedra introduced by Rademacher [6].)
Throughout the paper, all graphs are simple, even if this is not indicated separately in the statement.
Definition 1.1.
A Halin graph is a planar graph constructed by connecting the leaves of a tree into a cycle.
Definition 1.2.
A based planar graph is a planar graph that have a face that is adjacent to every other faces.
Remark 1.3.
Notice that a Halin graph is a based planar graph.
Definition 1.4.
[…]
Definition 1.5.
[…]
Remark 1.6.
For a based planar graph we never select the outer face in the cycle packing. This does not narrow the meaning of cycle packing, since we only delete one trivial case, but it makes the proofs technically simpler.
Definition 1.7.
[…]
Theorem 1.8.
[…]
Corollary 1.9.
[…]
2 Proof of Theorem 1.8
From now on, the outer face of a based planar graph is adjacent to every other faces.
Notation 2.1.
In the following proof good triangle is a triangle $xyz$, for which $d(x)=d(y)=3$ such that the vertices $x$ and $y$ are on the outer face. A good triangle for a vertex $u$ is a good triangle not containing $u$.
Claim 2.2.
If the degree of each vertex in a based planar graph $G$ is at least three and the degree of each vertex on the outer face is $3$, then for an arbitrary vertex $u$ on the outer face, the graph $G$ has a triangle $xyz$, for which $d(x)=d(y)=3$ such that the vertices $x$ and $y$ are on the outer face and $x\neq u,y\neq u$ and $z\neq u$.
Proof of Claim 2.2.
The proof is done by induction on the number of vertices.
If we omit the edges at the border of the outer face, we get a forest. If this forest is a tree, then let’s take a longest path in it. At least two neighbors of the penultimate vertex of this path will be leaves since its degree is at least three and apart from its neighbor on the path, it cannot have a neighbor that is not a leaf, because then we would not have started from the longest path. Two of these leaves will be adjacent on the outer face, so we have found a suitable triangle. We could do the same at the other end of the longest path. So, for any vertex, we find a suitable triangle that does not contain the given vertex.
Consider the case where the resulting forest consists of several components. Let us choose one of the trees as we like. Its leaves divide the boundary of the outer face into arcs. Let us choose another tree, its vertices are on one of the resulting arcs. We can assume that neither tree of the two has only two points. If this were the case, then we can find suitable trees on both arcs formed by the two-point tree, and we can choose them. The two end points of the arc should be the leaves $a$ and $b$, and the two extreme leaves of the second tree on the arc should be $c$ and $d$ such that $a$ is connected with $c$ and $b$ is connected with $d$. Let we delete the edges $ac$ and $bd$ and draw in the edges $ab$ and $cd$. Thus, we get two graphs for which we can use induction. If in both graphs we choose the point from which we want to find a disjoint triangle so that it is one of the endpoints of the new edge, then in both graphs we find a good triangle that does not use the new edge. Thus these good triangles were good triangles in the original graph too and they are disjoint, so the statement is holds. (It is not necessary, but it is sufficient to find two disjoint triangles whose vertices on the outer face have degree $3$.) ∎
Lemma 2.3.
If the degree of each vertex in a based planar graph $G$ is at least three, then for an arbitrary vertex $u$ on the outer face, the graph $G$ has a triangle $xyz$, for which $d(x)=d(y)=3$ such that the vertices $x$ and $y$ are on the outer face and $x\neq u,y\neq u$ and $z\neq u$.
Proof of Lemma 2.3.
Let $k$ be the number of vertices on the outer face whose degree is greater than three. The proof is done by induction according to $k$ and within it by induction according to the number of vertices. (So, if the value of $k$ is smaller for another graph, or the value of $k$ is the same, but the number of peaks is smaller, then according to the induction assumption, the statement is true for the other graph.) If $k=0$, then the statement is true beacuse of Claim 2.2.
If $k>0$, then let us consider a vertex $v$ on the outer face whose degree is greater than three. If the vertex $v$ is a cut vertex, then $G=G_{1}\cup G_{2}$ such that $G_{1}\cap G_{2}=\{v\}$ and $G_{1}\setminus\{v\}$ and $G_{2}\setminus\{v\}$ are disjoint. If the degree of the vertex $v$ in the graph $G_{1}$ or in the graph $G_{2}$ is at least three, then we are done. Since it can be assumed that, the degree of the vertex $v$ is at least three in the graph $G_{1}$. Then, for every vertex of the graph $G_{1}$ on the outer domain, there will be a good triangle due to induction (the graph $G_{1}$ has fewer vertices than the graph $G$ and the value of $k$ cannot increase with the operation) and any of these triangles will also be good for the vertices of the graph $G_{2}\setminus\{v\}$. If the degree of vertex $v$ in both graph $G_{1}$ and $G_{2}$ is two, then the neighbors of the vertex $v$ in the graph $G_{1}$ should be $e$ and $f$ and the neighbors of the vertex $v$ in the graph $G_{2}$ should be $g$ and $h$. Then we delete the vertex $v$ and drag the edges $eg$ and $fh$ into the graph. We can apply induction to the resulting graph. There will be a good triangle for each vertex and this triangle will be a good triangle in the original graph too, since we did not create a new triangle (and these triangles will be good for the vertex $v$ too).
Let us consider a vertex $v$ on the outer face whose degree is greater than three and what is not a cut vertex. Let’s split the vertex $v$ into a vertex $v_{1}$ and a vertex $v_{2}$. The two edges that are on the border of the outer face and run into the vertex $v$, after that let them run into the vertex $v_{1}$ and the other edges that run into the vertex $v$, after that let them run into the vertex $v_{2}$. This operation maintains that the degree of each vertex is at least three (the degree of the vertices $v_{1}$ and $v_{2}$ is at least three and the degrees of the other vertices do not change). It is preserved that the graph is a planar graph and that every face is adjacent to the outer face. The value of $k$ has decreased for the graph, so we can apply the induction assumption to it. The vertex $v_{1}$ will not be included in a triangle, if the vertex $v_{2}$ is included in a triangle, it corresponds to a triangle in the original graph in which the vertex $v$ is present instead of the vertex $v_{2}$.
The only problem is that the induction does not guarantee a good triangle for the vertex $v$, since for this we have to stipulate that the vertex $v_{2}$ is not included in the triangle, but the vertex $v_{2}$ is not on the border of the outer face. Assume by contradiction that every triangle that has two three-degrees vertices on the outer face contains the vertex $v$. There must be at least two such triangles, since for every vertex different from the vertex $v$ we proved by induction that there is a good triangle that does not contain that vertex. Let these two triangles be $vab$ and $vcd$ such that they are located in the order $vabcdv$ on the boundary of the outer face. The vertices $a$ and $d$ divide the boundary of the outer face into two arcs. Let’s take the one that contains the vertices $b$ and $c$. This arc, together with the vertices $a,d$, and $v$, defines a circle. Let us delete all the vertices that are in the circle and all internal points of the arc and join the vertices $a$ and $d$ with an edge. The resulting graph satisfies the conditions of the lemma: the degree of every vertex is at least three and every face is adjacent to the outer face. Furthermore, the number of vertices on the outer face whose degree is greater than three cannot increase. So we can use the induction assumption. In the resulting graph, we find a good triangle that does not contain $v$, this is also a suitable triangle in the original graph, so we have reached a contradiction. Since it cannot happen that we do not find a good triangle for the vertex $v$, the induction works and the statement of the lemma is true. ∎
Proof of Theorem 1.8.
The proof is done by induction on the number $n$ of vertices. If $n\leq 3$, then the theorem is trivially true. All subgraphs of a based planar graph is based planar graph.
If there is a vertex $u$ in the graph $G$ of which degree is smaller than $2$, then we leave one of these vertex Let the resulting graph be $G^{\prime}$. $\textrm{cp}(G)=\textrm{cp}(G^{\prime})$ and $\textrm{fvs}(G)=\textrm{fvs}(G^{\prime})$, since $u$ is not in any cycle. The statement is true for the graph $G^{\prime}$ by induction, so it is also true for the graph $G$.
If there is a triangle $xyz$ in the graph $G$ such that $d(x)=2$, then let $G^{\prime}$ be the graph obtained by omitting the vertices $y$ and $z$ from the graph $G$. $\textrm{cp}(G^{\prime})\leq\textrm{cp}(G)-1$ since let we take a cycle packing in the graph $G^{\prime}$, in the graph $G$ we can also add the triangle $xyz$ to this cycle packing. Thus for every cycle packing in the graph $G^{\prime}$ we find a larger cycle packing in the graph $G$, so the maximum size of a cycle packing of $G$ necessarily bigger than the maximum size of a cycle packing of $G^{\prime}$. Let be $\textrm{cp}(G)=k$ then $\textrm{cp}(G^{\prime})\leq k-1$, so $\textrm{fvs}(G^{\prime})\leq 2k-2$ by induction. Let be $F^{\prime}$ an arbitrary minimum feedback vertex set of the graph $G^{\prime}$, then $F=F^{\prime}\cup\{y,z\}$ is a minimum feedback vertex set of the graph $G$ because the difference between the graph $G\setminus F$ and the graph $G^{\prime}\setminus F^{\prime}$ is only an isolated vertex. Thus, $\textrm{fvs}(G)\leq\textrm{fvs}(G^{\prime})+2\leq 2k$.
If there is a vertex $u$ of degree two in the graph $G$, which is not in a triangle, then let $G^{\prime}$ be the graph obtained by omitting the vertex $u$ from the graph $G$ and connecting the two neighbors of the vertex $u$ with an edge. This graph is trivially still a based planar graph and all minimum feedback vertex set in this graph is also a minimum feedback vertex set in the original graph and vice versa. Furthermore, the same is true for cycle packings: they are also essentially identical in the two graphs. So $\textrm{cp}(G)=\textrm{cp}(G^{\prime})$ and $\textrm{fvs}(G)=\textrm{fvs}(G^{\prime})$, thus, by induction, we are done. (The statement was true for the graph $G^{\prime}$, so it is also true for the graph $G$.)
Now, we may assume that the degree of each vertex in the graph is at least three. Then, according to Lemma 2.3, there is a triangle $xyz$, for which $d(x)=d(y)=3$ such that the vertices $x$ and $y$ are on the outer face. Let $G^{\prime}$ be the graph obtained by deleting the vertices $y$ and $z$ from the graph $G$. $\textrm{cp}(G^{\prime})<\textrm{cp}(G)$, since the vertex $x$ cannot be included in any cycle in the graph $G^{\prime}$, so we can add the triangle $xyz$ to an arbitrary cycle packing of the graph $G^{\prime}$ and thus get a bigger cycle packing. Then let’s take a minimum feedback vertex set of the graph $G^{\prime}$, adding the vertices $y$ and $z$ to this, we get a minimum feedback vertex set of the graph $G$, whose size is at most $2\textrm{cp}(G^{\prime})+2\leq 2\cdot\textrm{cp}(G)$. ∎
We actually also proved the following stronger statement:
Corollary 2.4.
[…]
3 Concluding remarks
The proof takes advantage of the fact that there is a face that is adjacent to every other face. The question arises naturally whether the statement is true even if there is a cycle in the graph which has at least one edge in common with every face?
Problem 3.1.
[…]
Another similar approach might be the special case of hamiltonian graphs.
Problem 3.2.
[…]
Statements and declarations
The authors declare that they has no conflict of interest. Data sharing not applicable to this article as no datasets were generated or analysed during the current study.